OpenStudy (eamier):

prove that the derivative of log x base a equal to1/(a log x)

OpenStudy (anonymous):

\[y = log_ax\] \[->a^y=x\] the natural logarithm of both sides \[\ln a^y = \ln x\] rearrange \[y\ln a = \ln x\] implicit differentiation \[\ln ay^'\ = \frac{1}{x}\] \[y^'=\frac{1}{x\ln a}\]